OpenStudy (anonymous):

Find the reference angle for -690 degrees

OpenStudy (tkhunny):

-690 + 360 = ?? -690 + 360 + 360 = ?? Are we there yet?

OpenStudy (anonymous):

Is it 220?

OpenStudy (anonymous):

Or 140 degrees

OpenStudy (tkhunny):

No. Where did you get that? Note: It really is not helpful just to drop possible answers. Please show your efforts. Did you do what I suggested?

OpenStudy (tkhunny):

It's a Reference Angle. Doesn't it have to be between 0º and 90º (inclusive)?

OpenStudy (anonymous):

Yes. So -690 + 360 +360=30 degrees

OpenStudy (tkhunny):

That's it. Pop Quiz!!!! If you keep adding or subtracting 360º and you manage 135º when you decide to quit, what is the Reference Angle then?

OpenStudy (anonymous):

Would it be 135-90 degrees?

OpenStudy (tkhunny):

Yes, but that may be by accident. How about 175º? This will prove that you have it!